OpenStudy (winsteria):

MEDAL AND FAN What is the area of the triangle? A. 32 square centimeters B. 48 square centimeters C. 96 square centimeters D. 60 square centimeters

OpenStudy (anonymous):

pic

OpenStudy (trojanpoem):

1/2 * base * height = 8 * 0.5 * 12

OpenStudy (trojanpoem):

48

Nnesha (nnesha):

stop giving out direct answer

Nnesha (nnesha):

all of you

OpenStudy (trojanpoem):

0.5 base * height ? I don't think it need more explanation.

OpenStudy (winsteria):

It wasn't exactly direct he gave out the formula to do it and thank you.

OpenStudy (anonymous):

Ha ha ha ha.. @TrojanPoem

OpenStudy (anonymous):

\[\large \color{green}{Area_{\triangle} = \frac{1}{2} \times Base \times Height}\]

Nnesha (nnesha):

first of all formulas suppose to be on notes every question suppose to have a little bit explanation :) just give him/her formula and let them use calculator its fun to play with calculator like iphone

Nnesha (nnesha):

:-)

OpenStudy (trojanpoem):

@eta , you rocks ++.

OpenStudy (anonymous):

Calculator hinders your real mental ability.. Stay away from it wherever possible..!! :)

OpenStudy (anonymous):

your 'hindi' is too good.. :P

OpenStudy (trojanpoem):

@eta, respect ++, Calculator = bye mathematics